# Self-adjoint traces on the Pedersen ideal of $\mathrm{C}^{\ast}$-algebras

### v1

@@ 1. Introduction
The invariant – which we denote by $T_{\mathbb{R}}(A)$ – is the space of self-adjoint traces inside $\mathrm{Ped}(A)^{\ast}$ equipped with the weak-topology. This turns out to be the span of $T_{+}(A)$ (Corollary 3.7), so it might be hard to see how this contains any new information compared to $T_{+}(A)$. However, the following observation – that the related Trace Question for $T_{\mathbb{R}}(A)$ has a positive answer – motivates considering $T_{\mathbb{R}}(A)$ instead of $T_{+}(A)$.
@@
In order to systematically study $T_{\mathbb{R}}(A)$ we introduce a topology on $\mathrm{Ped}(A)$ as follows. Since $\mathrm{Ped}(A)$ is the directed union of all its $\mathrm{C}^{\ast}$-subalgebras (Remark 2.3), we equip $\mathrm{Ped}(A)$ with the induced locally convex inductive limit topology. The continuous dual space $\mathrm{Ped}(A)^{\ast}$ thus consists exactly of the linear functionals on $\mathrm{Ped}(A)$ which are bounded when restricted to any $\mathrm{C}^{\ast}$-subalgebra of $\mathrm{Ped}(A)$. In particular, all positive linear functionals on $\mathrm{Ped}(A)$ are continuous. Formally, we define $T_{\mathbb{R}}(A)\subseteq\mathrm{Ped}(A)^{\ast}$ as the space of continuous self-adjoint traces. This is a real topological vector space which is ordered by the cone $T_{+}(A)$ of positive traces. Moreover, by utilising Jordan decompositions of linear functionals we give a different proof (Corollary 3.8) of a result by Pedersen [Ped69, Theorem 3.1] that $T_{\mathbb{R}}(A)$ is a vector lattice. [^1] In the case where $A$ is unital, this gives an easy proof that the set of tracial states is a Choquet simplex.
@@
Let $A$ and $B$ be $\mathrm{C}^{\ast}$-algebras. Then $T_{\mathbb{R}}(A)$ is canonically isomorphic to the space of compact-open-continuous linear functionals on $T_{\mathbb{R}}(A)^{\ast}$. This induces a contravariant categorical duality, namely a bijective correspondence between weak-continuous positive linear maps $T_{\mathbb{R}}(B)\to T_{\mathbb{R}}(A)$ and compact-open-continuous positive linear maps $T_{\mathbb{R}}(A)^{\ast}\to T_{\mathbb{R}}(B)^{\ast}$.
@@
For any locally compact Hausdorff space $X$, $T_{\mathbb{R}}(C_{0}(X))$ is canonically identified [^2] In fact, this is how real Radon measures on $X$ are defined in the Bourbaki book [Bou52, Chapter III]. with the space of real Radon measures $\mathrm{Rad}_{\mathbb{R}}(X)$ on $X$ with the weak-topology coming from $C_{c}(X)=\mathrm{Ped}(C_{0}(X))$. We study traces more generally on the $\mathrm{C}^{\ast}$-algebra of an étale groupoid $\mathcal{G}$, which have been considered for instance in [Nes13, LR19, AS21, Chr23, LZ24]; we are especially interested in the entire structure of the trace space, notably its topology. Consider the set $\mathrm{Rad}_{\mathbb{R}}^{\mathcal{G}}(\mathcal{G}^{0})$ of $\mathcal{G}$-invariant real Radon measures on the unit space $\mathcal{G}^{0}$ equipped with the weak-topology coming from $C_{c}(\mathcal{G}^{0})$. We show (Corollary 7.5) that for a twisted étale groupoid $(\mathcal{G},\mathcal{L})$ (not necessarily Hausdorff), every $\mu\in\mathrm{Rad}_{\mathbb{R}}^{\mathcal{G}}(\mathcal{G}^{0})$ induces a trace $\tau_{\mu}\in T_{\mathbb{R}}(\mathrm{C}^{\ast}_{\lambda}(\mathcal{G},\mathcal{L}))$ which is uniquely determined by
$\tau_{\mu}(f)=\int_{\mathcal{G}^{0}}f|_{\mathcal{G}^{0}}\mathrm{d}\mu,\qquad f\in\Gamma_{c}(\mathcal{G},\mathcal{L}).$
Moreover, when $\mathcal{G}$ is Hausdorff the weak-topology on this set of traces coming from $\mathrm{Ped}(\mathrm{C}^{\ast}_{\lambda}(\mathcal{G},\mathcal{L}))$ agrees with the one coming from $C_{c}(\mathcal{G}^{0})$ (Lemma 7.8). In the case where $\mathcal{G}$ is principal, this determines the entire space of traces (Corollary 7.9).
@@ Acknowledgement
The authors would like to thank Leonel Robert for an email correspondence on the Trace Question. The first named author would also like to thank José Carrión, Jorge Castillejos, Sam Evington, Chris Schafhauser, Aaron Tikuisis, and Stuart White for enlightening conversations throughout the years about bounded traces.
@@ 2. Continuous traces on the Pedersen ideal
We will be considering the Pedersen ideal with a certain locally convex structure which in general is stronger than the usual norm topology. The starting point is the following important result of Pedersen.
@@
It suffices to consider the case $n=2$, so let $B_{0}=\mathrm{C}^{\ast}(D_{1},D_{2})$ and let $B$ be the hereditary $\mathrm{C}^{\ast}$-subalgebra generated by $B_{0}$. Fix $x\in B$. By Cohen’s factorisation theorem (see [BO08, Theorem 4.6.4] for an easy proof in the case that we need it for) we have $B=B_{0}A$ so we may write $x=by$ for a $b\in B_{0}$ and $y\in A$. Pick separable $\mathrm{C}^{\ast}$-subalgebras $C_{j}\subseteq D_{j}$ for $j=1,2$ such that $b\in\mathrm{C}^{\ast}(C_{1},C_{2})$. As $C_{j}$ is separable and thus $\sigma$-unital we let $h_{j}\in C_{j}$ be strictly positive. Then $h_{1},h_{2}\in\mathrm{Ped}(A)$ and by Lemma 2.1 the hereditary $\mathrm{C}^{\ast}$-subalgebra $C$ generated by $\{h_{1},h_{2}\}$ is contained in $\mathrm{Ped}(A)$. Since $b\in\mathrm{C}^{\ast}(C_{1},C_{2})\subseteq C\subseteq\mathrm{Ped}(A)$, and since $\mathrm{Ped}(A)$ is an ideal, it follows that $x=by\in\mathrm{Ped}(A)$. ∎
@@
This way of defining continuous functionals on $\mathrm{Ped}(A)$ is similar to (and inspired by) how one defines real or complex Radon measures on locally compact spaces $X$. For instance, as done in the Bourbaki book [Bou52, Chapter III], this is done by identifying complex Radon measures on $X$ with linear functionals on $C_{c}(X)=\mathrm{Ped}(C_{0}(X))$ (the space of compactly supported functions) which are continuous when $C_{c}(X)$ is equipped with the the same locally convex inductive limit topology as described above.
@@
We often consider $T_{\mathbb{R}}(A)$ as a space of functionals on the real vector space $\mathrm{Ped}(A)_{\mathrm{sa}}$ of self-adjoint elements in $\mathrm{Ped}(A)$. Each $\tau\in T_{\mathbb{R}}(A)$ restricts to a functional $\tau\colon\mathrm{Ped}(A)_{\mathrm{sa}}\to\mathbb{R}$ which is tracial in the sense that $\tau(x^{\ast}x)=\tau(xx^{\ast})$ for all $x\in\mathrm{Ped}(A)$. It is straightforward to see that, conversely, any linear functional $\tau\colon\mathrm{Ped}(A)_{\mathrm{sa}}\to\mathbb{R}$ with this tracial property extends uniquely to a self-adjoint tracial functional $\tilde{\tau}$ on $\mathrm{Ped}(A)$. The extension $\tilde{\tau}$ is continuous if and only if $\tau\colon T_{\mathbb{R}}(A)_{\mathrm{sa}}\to\mathbb{R}$ is continuous, so we may identify the self-adjoint traces as a subspace $T_{\mathbb{R}}(A)\subseteq(\mathrm{Ped}(A)_{\mathrm{sa}})^{*}$, where its topology agrees with the weak topology. This will be used often without reference.
@@ 3. Jordan decomposition
Proposition 3.1(c) gives an elementary proof (assuming the existence of Jordan decompositions of bounded self-adjoint functionals) that the set of bounded self-adjoint tracial functionals on $\mathrm{C}^{\ast}$-algebras form a vector lattice with join $\tau_{1}\vee\tau_{2}=(\tau_{1}-\tau_{2})^{+}+\tau_{2}$ and meet $\tau_{1}\wedge\tau_{2}=\tau_{1}-(\tau_{1}-\tau_{2})^{+}$. In particular, the set of tracial states on a unital $\mathrm{C}^{\ast}$-algebra is a Choquet simplex.
@@
If $D$ is a $\mathrm{C}^{\ast}$-algebra and $(\tau_{i})$ is a bounded net of self-adjoint tracial functionals converging weak to $\tau$, and if $(\tau_{i}^{+},\tau_{i}^{-})$ and $(\tau^{+},\tau^{-})$ are the respective Jordan decompositions, then it is in general not true that $\tau_{i}^{+}\to\tau^{+}$ and $\tau_{i}^{-}\to\tau^{-}$ (even when we know that $(\tau_{i}^{+})$ and $(\tau_{i}^{-})$ both converge to some tracial functionals).
A counter example could be $D=C([-1,1])$ and $\tau_{n}=\mathrm{ev}_{1/n}-\mathrm{ev}_{-1/n}$ (where $\mathrm{ev}_{t}$ is evaluation at $t$) which converges weak to 00. However, $\tau_{n}^{\pm}=\mathrm{ev}_{\pm 1/n}$ which both converge weak to $\mathrm{ev}_{0}$.
@@ 4. The dual trace space
Let $A$ be a $\mathrm{C}^{\ast}$-algebra and let $\rho\colon\mathrm{Ped}(A)\to[0,\infty)$ be a seminorm such that the restriction $\rho|_{D}$ is bounded for every $\mathrm{C}^{\ast}$-subalgebra $D\subseteq\mathrm{Ped}(A)$. Define the compact set
@@
Then for every $f\in T_{\mathbb{R}}(A)^{\ast}$ and every $\epsilon>0$ there exists an $a\in\mathrm{Ped}(A)_{\mathrm{sa}}$ such that $\hat{a}=f$ and $\rho(a)\leq\|f\|_{K_{\rho}}+\epsilon$.
@@
Suppose $A$ is a $\mathrm{C}^{\ast}$-algebra such that all non-zero traces are unbounded. From Corollary 4.5 it follows that every $f\in T_{\mathbb{R}}(A)^{\ast}$ can be realised as $f=\hat{a}$ for a self-adjoint $a\in\mathrm{Ped}(A)_{\mathrm{sa}}$ with arbitrarily small norm. This is easy to see if $A$ is stable, but was somewhat surprising to the authors in general.
@@
We will use Theorem 4.3 several times to identify elemenents of $T_{\mathbb{R}}(D)^{\ast}$ with self-adjoint elements in $\mathrm{Ped}(D)$.
@@
and similarly $\tau(\phi(a))\leq\tau(\psi(a))$. As $T_{\mathbb{R}}(B)$ and $\mathrm{Ped}(A)$ are spanned by $T_{+}(B)$ and $\mathrm{Ped}(A)_{+}$ respectively, the result follows.
@@ 5. Extending traces from subalgebras
(b): Let $a_{1},\dots,a_{n}\in\mathrm{Ped}(A)_{0}$ such that $a\leq\sum_{j}a_{j}$. By [Ped79, Proposition 1.4.10] there are $v_{1},\dots,v_{n}\in A$ such that $v_{j}v_{j}^{\ast}\leq a_{j}$ (and in particular, $v_{j}v_{j}^{\ast}\in\mathrm{Ped}(A)_{0}$), and $a=\sum_{j=1}^{n}v_{j}^{\ast}v_{j}$. Let $e_{j}\in\mathrm{Ped}(A)_{+}$ so that $e_{j}v_{j}=v_{j}$. By part (a) we may find $x_{j,k}\in X$ (finitely many) and $z_{j,k}\in A$ such that $\|e_{j}-\sum_{k}z_{j,k}^{\ast}x_{j,k}^{\ast}x_{j,k}z_{j,k}\|<1/2$. By [KR02, Lemma 2.2] find $d_{j,k}\in A$ so that $\sum_{k}d_{j,k}^{\ast}z_{j,k}^{\ast}x_{j,k}^{\ast}x_{j,k}z_{j,k}d_{j,k}=e_{j}^{\prime}:=2(e_{j}-\tfrac{1}{2})_{+}$. Then $e_{j}^{\prime}v_{j}=v_{j}$. [^4] In fact, if $f\in C_{0}((0,\|e_{j}\|\,])$ then $f(e_{j})v_{j}=f(1)v_{j}$, since this can be verified when $f$ is a polynomial with no constant term. Letting $c_{j,k}:=z_{j,k}d_{j,k}v_{j}$ we get
$\sum_{j,k}c_{j,k}^{\ast}x_{j,k}^{\ast}x_{j,k}c_{j,k}=\sum_{j}v_{j}^{\ast}e_{j}^{\prime}v_{j}=a.\qed$
The following proposition gives a sufficient condition for when traces on $A$ can be studied through a $\ast$-subalgebra $A_{0}\subseteq\mathrm{Ped}(A)$. We have stated the result somewhat generally (and technically) so that it can hopefully be useful in multiple contexts, such as when $A_{0}$ is dense or a full hereditary $\mathrm{C}^{\ast}$-subalgebra.
@@
Let $A$ and $B$ be $\mathrm{C}^{\ast}$-algebras with traces $\tau_{A}\in T_{\mathbb{R}}(A)$ and $\tau_{B}\in T_{\mathbb{R}}(B)$. Then there is a unique trace $\tau_{A}\otimes\tau_{B}\in T_{\mathbb{R}}(A\otimes B)$ which satisfies $(\tau_{A}\otimes\tau_{B})(a\otimes b)=\tau_{A}(a)\tau_{B}(b)$ for each $a\in\mathrm{Ped}(A)$ and $b\in\mathrm{Ped}(B)$ and is positive if both $\tau_{A}$ and $\tau_{B}$ are positive. Moreover, for any $\tau_{B}\in T_{\mathbb{R}}(B)$ the assignment $\tau_{A}\mapsto\tau_{A}\otimes\tau_{B}\colon T_{\mathbb{R}}(A)\to T_{\mathbb{R}}(A\otimes B)$ is continuous. If $\tau_{B}$ is the unique non-zero positive trace in $T_{\mathbb{R}}(B)$ up to scaling then this is an isomorphism of ordered topological vector spaces.

### v2

@@ Abstract.
Key words and phrases:
Trace space, non-unital $\mathrm{C}^{\ast}$-algebra, Pedersen ideal
@@ 1. Introduction
The invariant – which we denote by $T_{\mathbb{R}}(A)$ – is the subspace of $\mathrm{Ped}(A)^{\ast}$ consisting of self-adjoint traces equipped with the weak-topology. This turns out to be the span of $T_{+}(A)$ (Corollary 3.7), so it might be hard to see how this contains any new information compared to $T_{+}(A)$. However, the following observation – that the related Trace Question for $T_{\mathbb{R}}(A)$ has a positive answer – motivates considering $T_{\mathbb{R}}(A)$ instead of $T_{+}(A)$.
@@
In order to systematically study $T_{\mathbb{R}}(A)$ we introduce a topology on $\mathrm{Ped}(A)$ as follows. Since $\mathrm{Ped}(A)$ is the directed union of all its $\mathrm{C}^{\ast}$-subalgebras (Remark 2.3), we equip $\mathrm{Ped}(A)$ with the induced locally convex inductive limit topology. The continuous dual space $\mathrm{Ped}(A)^{\ast}$ thus consists exactly of the linear functionals on $\mathrm{Ped}(A)$ which are bounded when restricted to any $\mathrm{C}^{\ast}$-subalgebra of $\mathrm{Ped}(A)$. In particular, all positive linear functionals on $\mathrm{Ped}(A)$ are continuous. Formally, we define $T_{\mathbb{R}}(A)\subseteq\mathrm{Ped}(A)^{\ast}$ as the space of continuous self-adjoint traces. This is a real topological vector space which is ordered by the cone $T_{+}(A)$ of positive traces. Moreover, by utilising Jordan decompositions of linear functionals we give a different proof (Corollary 3.8) of a result by Pedersen [Ped69, Theorem 3.1] that $T_{\mathbb{R}}(A)$ is a vector lattice. [^1] In the case where $A$ is unital, this gives an easy proof that the set of tracial states is a Choquet simplex. See [BR24] for another recent short proof of this result.
@@
Let $A$ and $B$ be $\mathrm{C}^{\ast}$-algebras. Then $T_{\mathbb{R}}(A)$ is canonically isomorphic to the space of compact-open-continuous linear functionals on $T_{\mathbb{R}}(A)^{\ast}$. This induces a contravariant categorical duality, that is, a bijective correspondence between weak-continuous positive linear maps $T_{\mathbb{R}}(B)\to T_{\mathbb{R}}(A)$ and compact-open-continuous positive linear maps $T_{\mathbb{R}}(A)^{\ast}\to T_{\mathbb{R}}(B)^{\ast}$.
@@
For any locally compact Hausdorff space $X$, the real trace space $T_{\mathbb{R}}(C_{0}(X))$ is frequently identified [^2] In fact, this is how real Radon measures on $X$ are defined in the Bourbaki book [Bou52, Chapter III]. with the space of real Radon measures $\mathrm{Rad}_{\mathbb{R}}(X)$ on $X$ with the weak-topology coming from $C_{c}(X)=\mathrm{Ped}(C_{0}(X))$, via the map sending a measure $\mu$ to the trace $f\mapsto\int_{X}f\mathrm{d}\mu$. We study traces more generally on the $\mathrm{C}^{\ast}$-algebra of an étale groupoid $\mathcal{G}$, which have been considered for instance in [Nes13, LR19, AS21, NS22, Chr23, LZ24]; we are especially interested in the entire structure of the trace space, notably its topology. Consider the set $\mathrm{Rad}_{\mathbb{R}}^{\mathcal{G}}(\mathcal{G}^{0})$ of $\mathcal{G}$-invariant real Radon measures on the unit space $\mathcal{G}^{0}$ equipped with the weak-topology coming from $C_{c}(\mathcal{G}^{0})$. We show (Corollary 7.5) that for a twisted étale groupoid $(\mathcal{G},\mathcal{L})$ (not necessarily Hausdorff), every $\mu\in\mathrm{Rad}_{\mathbb{R}}^{\mathcal{G}}(\mathcal{G}^{0})$ induces a trace $\tau_{\mu}\in T_{\mathbb{R}}(\mathrm{C}^{\ast}_{\lambda}(\mathcal{G},\mathcal{L}))$ which is uniquely determined by
$\tau_{\mu}(f)=\int_{\mathcal{G}^{0}}f|_{\mathcal{G}^{0}}\mathrm{d}\mu$
for compactly supported continuous [^3] If $\mathcal{G}$ is not Hausdorff these need not actually be continuous, see Section 7. sections $f\in\Gamma_{c}(\mathcal{G},\mathcal{L})$ of the line bundle $\mathcal{L}\to\mathcal{G}$. Moreover, when $\mathcal{G}$ is Hausdorff the weak-topology on this set of traces coming from $\mathrm{Ped}(\mathrm{C}^{\ast}_{\lambda}(\mathcal{G},\mathcal{L}))$ agrees with the one coming from $C_{c}(\mathcal{G}^{0})$ (Lemma 7.8). In the case where $\mathcal{G}$ is principal, this determines the entire space of traces (Corollary 7.9).
@@ Acknowledgement
The authors would like to thank Leonel Robert for an email correspondence on the Trace Question. The first named author would also like to thank José Carrión, Jorge Castillejos, Sam Evington, Chris Schafhauser, Aaron Tikuisis, and Stuart White for enlightening conversations throughout the years about bounded traces. The authors would also like to thank the referees for their very helpful comments.
@@ 2. Continuous traces on the Pedersen ideal
We will be considering the Pedersen ideal with a locally convex structure which in general is stronger than the usual norm topology. The starting point is the following important result of Pedersen.
@@
Since we may iteratively construct the final hereditary $\mathrm{C}^{\ast}$-subalgebra by adding in one $D_{k}$ at a time, it suffices to consider the case $n=2$. Let $B_{0}=\mathrm{C}^{\ast}(D_{1},D_{2})$ and let $B$ be the hereditary $\mathrm{C}^{\ast}$-subalgebra generated by $B_{0}$. Fix $x\in B$. By Cohen’s factorisation theorem (see [BO08, Theorem 4.6.4] for an easy proof in the case that we need it for) we have $B=B_{0}A$ so we may write $x=by$ for a $b\in B_{0}$ and $y\in A$. Pick separable $\mathrm{C}^{\ast}$-subalgebras $C_{j}\subseteq D_{j}$ for $j=1,2$ such that $b\in\mathrm{C}^{\ast}(C_{1},C_{2})$. As $C_{j}$ is separable and thus $\sigma$-unital we let $h_{j}\in C_{j}$ be strictly positive. Then $h_{1},h_{2}\in\mathrm{Ped}(A)$ and by Lemma 2.1 the hereditary $\mathrm{C}^{\ast}$-subalgebra $C$ generated by $\{h_{1},h_{2}\}$ is contained in $\mathrm{Ped}(A)$. Since $b\in\mathrm{C}^{\ast}(C_{1},C_{2})\subseteq C\subseteq\mathrm{Ped}(A)$, and since $\mathrm{Ped}(A)$ is an ideal, it follows that $x=by\in\mathrm{Ped}(A)$. ∎
@@
This way of defining continuous functionals on $\mathrm{Ped}(A)$ is similar to (and inspired by) the way in which one defines real or complex Radon measures on locally compact spaces $X$. For instance, as done in the Bourbaki book [Bou52, Chapter III], this is done by identifying complex Radon measures on $X$ with linear functionals on $C_{c}(X)=\mathrm{Ped}(C_{0}(X))$ (the space of compactly supported functions) which are continuous when $C_{c}(X)$ is equipped with the same locally convex inductive limit topology as described above.
@@
We often consider $T_{\mathbb{R}}(A)$ as a space of functionals on the real vector space $\mathrm{Ped}(A)_{\mathrm{sa}}$ of self-adjoint elements in $\mathrm{Ped}(A)$. Each $\tau\in T_{\mathbb{R}}(A)$ restricts to a functional $\tau\colon\mathrm{Ped}(A)_{\mathrm{sa}}\to\mathbb{R}$ which is tracial in the sense that $\tau(x^{\ast}x)=\tau(xx^{\ast})$ for all $x\in\mathrm{Ped}(A)$. It is straightforward to see that, conversely, any linear functional $\tau\colon\mathrm{Ped}(A)_{\mathrm{sa}}\to\mathbb{R}$ with this tracial property extends uniquely to a self-adjoint tracial functional $\tilde{\tau}$ on $\mathrm{Ped}(A)$. The extension $\tilde{\tau}$ is continuous if and only if $\tau\colon\mathrm{Ped}(A)_{\mathrm{sa}}\to\mathbb{R}$ is continuous, so we may identify the self-adjoint traces as a subspace $T_{\mathbb{R}}(A)\subseteq(\mathrm{Ped}(A)_{\mathrm{sa}})^{*}$, where its topology agrees with the weak topology. This will be used often without reference.
@@ 3. Jordan decomposition
Proposition 3.1(c) gives an elementary proof (assuming the existence of Jordan decompositions of bounded self-adjoint functionals) that the set of bounded self-adjoint tracial functionals on a $\mathrm{C}^{\ast}$-algebra forms a vector lattice with join $\tau_{1}\vee\tau_{2}=(\tau_{1}-\tau_{2})^{+}+\tau_{2}$ and meet $\tau_{1}\wedge\tau_{2}=\tau_{1}-(\tau_{1}-\tau_{2})^{+}$. In particular, the set of tracial states on a unital $\mathrm{C}^{\ast}$-algebra is a Choquet simplex.
@@
If $D$ is a $\mathrm{C}^{\ast}$-algebra and $(\tau_{i})$ is a bounded net of self-adjoint tracial functionals converging weak to $\tau$, and if $(\tau_{i}^{+},\tau_{i}^{-})$ and $(\tau^{+},\tau^{-})$ are the respective Jordan decompositions, then it is in general not true that $\tau_{i}^{+}\to\tau^{+}$ and $\tau_{i}^{-}\to\tau^{-}$ (even when we know that $(\tau_{i}^{+})$ and $(\tau_{i}^{-})$ both converge to tracial functionals).
For example, let $D=C([-1,1])$ and $\tau_{n}=\mathrm{ev}_{1/n}-\mathrm{ev}_{-1/n}$ (where $\mathrm{ev}_{t}$ is evaluation at $t$). Then $\tau_{n}\xrightarrow{w^{\ast}}0$ but $\tau_{n}^{\pm}=\mathrm{ev}_{\pm 1/n}\xrightarrow{w^{\ast}}\mathrm{ev}_{0}\neq 0$.
@@ 4. The dual trace space
Let $A$ be a $\mathrm{C}^{\ast}$-algebra and let $\rho\colon\mathrm{Ped}(A)\to[0,\infty)$ be a seminorm such that the restriction $\rho|_{D}$ is bounded for every $\mathrm{C}^{\ast}$-subalgebra $D\subseteq\mathrm{Ped}(A)$. Let
@@
Then $K_{\rho}$ is compact, and for every $f\in T_{\mathbb{R}}(A)^{\ast}$ and every $\epsilon>0$ there exists an $a\in\mathrm{Ped}(A)_{\mathrm{sa}}$ such that $\hat{a}=f$ and $\rho(a)\leq\|f\|_{K_{\rho}}+\epsilon$.
@@
Suppose $A$ is a $\mathrm{C}^{\ast}$-algebra whose non-zero traces are all unbounded. From Corollary 4.5 it follows that every $f\in T_{\mathbb{R}}(A)^{\ast}$ can be realised as $f=\hat{a}$ for a self-adjoint $a\in\mathrm{Ped}(A)_{\mathrm{sa}}$ with arbitrarily small norm. This is easy to see if $A$ is stable, but was somewhat surprising to the authors in general.
@@
We will use Theorem 4.3 several times to identify elements of $T_{\mathbb{R}}(D)^{\ast}$ with self-adjoint elements in $\mathrm{Ped}(D)$.
@@
and similarly $\tau(\phi(a))\leq\tau(\psi(a))$. As $T_{\mathbb{R}}(B)$ and $\mathrm{Ped}(A)$ are spanned by $T_{+}(B)$ and $\mathrm{Ped}(A)_{+}$ respectively, the equality $T_{\mathbb{R}}(\phi)=T_{\mathbb{R}}(\psi)$ follows.
@@ 5. Extending traces from subalgebras
(b): Let $a_{1},\dots,a_{n}\in\mathrm{Ped}(A)_{0}$ such that $a\leq\sum_{j}a_{j}$. By [Ped79, Proposition 1.4.10] there are $v_{1},\dots,v_{n}\in A$ such that $v_{j}v_{j}^{\ast}\leq a_{j}$ (and in particular, $v_{j}v_{j}^{\ast}\in\mathrm{Ped}(A)_{0}$), and $a=\sum_{j=1}^{n}v_{j}^{\ast}v_{j}$. Let $e_{j}\in\mathrm{Ped}(A)_{+}$ so that $e_{j}v_{j}=v_{j}$. By part (a) we may find $x_{j,k}\in X$ (finitely many) and $z_{j,k}\in A$ such that $\|e_{j}-\sum_{k}z_{j,k}^{\ast}x_{j,k}^{\ast}x_{j,k}z_{j,k}\|<1/2$. By [KR02, Lemma 2.2] find $d_{j}\in A$ so that $\sum_{k}d_{j}^{\ast}z_{j,k}^{\ast}x_{j,k}^{\ast}x_{j,k}z_{j,k}d_{j}=e_{j}^{\prime}:=(e_{j}-\tfrac{1}{2})_{+}$. Then $e_{j}^{\prime}v_{j}=\tfrac{1}{2}v_{j}$. [^5] In fact, if $f\in C_{0}((0,\|e_{j}\|\,])$ then $f(e_{j})v_{j}=f(1)v_{j}$, since this can be verified when $f$ is a polynomial with no constant term. Letting $c_{j,k}:=\sqrt{2}z_{j,k}d_{j}v_{j}$ we get
$\sum_{j,k}c_{j,k}^{\ast}x_{j,k}^{\ast}x_{j,k}c_{j,k}=2\sum_{j}v_{j}^{\ast}e_{j}^{\prime}v_{j}=a.\qed$
The following proposition gives a sufficient condition under which traces on $A$ can be studied through a $\ast$-subalgebra $A_{0}\subseteq\mathrm{Ped}(A)$. We have stated the result somewhat generally (and technically) so that it can hopefully be useful in multiple contexts, such as when $A_{0}$ is dense or a full hereditary $\mathrm{C}^{\ast}$-subalgebra.
@@
Let $A$ and $B$ be $\mathrm{C}^{\ast}$-algebras with traces $\tau_{A}\in T_{\mathbb{R}}(A)$ and $\tau_{B}\in T_{\mathbb{R}}(B)$. Then there is a unique trace $\tau_{A}\otimes\tau_{B}\in T_{\mathbb{R}}(A\otimes B)$ that satisfies $(\tau_{A}\otimes\tau_{B})(a\otimes b)=\tau_{A}(a)\tau_{B}(b)$ for each $a\in\mathrm{Ped}(A)$ and $b\in\mathrm{Ped}(B)$ and is positive if both $\tau_{A}$ and $\tau_{B}$ are positive. Moreover, for any $\tau_{B}\in T_{\mathbb{R}}(B)$ the assignment $\tau_{A}\mapsto\tau_{A}\otimes\tau_{B}\colon T_{\mathbb{R}}(A)\to T_{\mathbb{R}}(A\otimes B)$ is continuous. If $\tau_{B}$ is the unique non-zero positive trace in $T_{\mathbb{R}}(B)$ up to scaling then this is an isomorphism of ordered topological vector spaces.